# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 4.2.5. Workflows have higher costs than zero-shot prompting
Figure shows two flow charts, one for crowdsourcing Mechanical Novel and one for LLM chain Mechanical Novel. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) additional verification steps in the reflect part of the workflow, 2) an automated length verification step in the revise part of the workflow 3) extra iterations through the scenes of the text as well as with different seeded texts.
@@ 4.3. Case study 3: Short story generation
The final case study concerns story generation given a short text prompt. This task lies on the open-ended creativity side of the output quality spectrum. Short stories have a wide design space of possible plots, characters, messages, imagery, and lessons. In this task, user revision of the story involves higher level control of different plot elements, with the ability to see counterfactuals (e.g., if a story element is included or not). We implement a workflow that, compared to zero-shot approaches, enables this exploration more precisely, while also creating stories that are preferred by crowdsourced raters for imagery, originality, and style.
@@ 4.3.1. Workflow design
We adapt the Mechanical Novel (MN) (Kim et al., 2017) workflow for this task. This workflow involves three main steps (Figure 10). The first step “initializes” a story with multiple scenes based on a given story prompt (e.g., “Malcolm finds himself alone in a runaway hot air balloon and accidentally travels to a city in the sky.”). Then, the “reflect” and “revise” steps are iterated five times in a loop. In the “reflect” step, workers consider what changes might helpfully adjust the story. In the “revise” step, workers implement those changes. Each step has subtasks for parallel generation of suggestions and edits, as well as voting for the best generation.
We adapted the Crowdsourced MN to MN-LLM by incorporating fine-grained user revision with manipulatable outputs. At the conclusion of the workflow, instead of receiving just one or more separate short stories, the user can interact with different combinations of edits to compare stories with different quality combinations and compare counterfactuals on editing decisions. To enable these outputs, we run the “revise” stage on texts of all combinations of suggestions from the “reflect” stage.
Beyond introducing control of story elements, the largest challenge was managing token context limitations. Getting diversity for creativity with iterative improvement on texts often increases the text length (Little et al., 2010a). To account for context length, we generate stories with four scenes rather than six and iterate through scenes rather than informing workers to select scenes. At the beginning of the “revise” step, we add a generation subtask that summarizes the current story to provide abbreviated context in the worker instructions. We included a hard-coded validation step to disqualify edits that were more than times the input text length. Unfortunately, this disqualified many highly descriptive texts and sometimes meant that no edits were made because all suggestions were too long.
Our initial pilots generated editing suggestions that were very generic (e.g., “Add more detail to the story”) and were repetitive across iteration rounds. To improve the diversity of suggestion outputs, we tried changing the worker instructions to be contrastive by providing previously used “banned” suggestions. However, the outputs anecdotally seemed more likely to follow those of prior rounds after this change. We hypothesize that the length of the context and request to do two tasks in one (generate a suggestion and avoid the prior suggestions) led to a lack of diversity. Instead, we add several verification steps to promote suggestion diversity and quality. This subtask granularity worked better. In these evaluate subtasks, we prompted the LLM to ensure that 1) the suggestion did not overlap with prior suggestions or two seeded generic suggestions (“Generic recommendation to introduce conflict.” and “Generic recommendation to add detail.”), 2) the suggestion has not already been implemented in the story, and 3) that the suggestion is worded as a suggestion, not new text for the story.
4.3.2. Evaluation setup
We generate one story in the zero-shot and workflow conditions for each of the five short story prompts in the Mechanical Novel paper (Kim et al., 2017), following their evaluation procedure.
4.3.3. Workflows enable control of plot elements
As shown in Table 3, MN-LLM generates more, longer, and higher quality suggestion combinations than the zero-shot baseline. MN-LLM does not always generate the 33 story options afforded by the initial story and combinations of 5 suggestions, as sometimes no generated suggestions passed our verification steps, or all generated fixes were too long. However, it still generates more than options on average, in comparison to zero-shot’s average of .
@@ 5. Discussion
The approach of decomposing a task into a series of subtasks is conceptually the same for both crowdsourcing workflows and LLM chains. Through overlap in the literature and our case study observations, we find that the high-level concepts of the design space apply to both, with the exception of the choice of workers. However, LLMs and people have different abilities, strengths, and weaknesses, so we may need different approaches for tactic designs to effectively implement the same strategies that achieve desired objectives. Thus, objectives, strategies, and some types of tactics transfer directly from crowdsourcing to LLMs. However, tactical design decisions can vary significantly, and could benefit from improved development and testing processes.
We first discuss transferrable approaches to user revision and creative work from crowdsourcing that could inspire a broader range of LLM chains. We then delve into findings from the literature and our case studies on how LLMs chain designs can differ from workflows powered by crowdworkers. Finally, the process itself of designing LLM chains is different because the running of LLMs is cheaper and faster. Using our experience from the case studies, we identify future directions to support chain design and prevent design tunneling. Finally, we summarize several concurrent papers investigating the intersection of LLMs and crowdsourcing.
5.1. What transfers from crowdsourcing
From the literature, we see that many tasks and all categories of the design space appear for both crowdsourcing workflows and LLM chains. The quality-cost tradeoff also appears in both literatures, although LLMs are less costly to query (Terwiesch and Ulrich, [n. d.]). Through our case studies, we find that we can also transfer new tasks (e.g., shortening and taxonomy creation) and underexplored strategies (e.g., direct manipulation, a subset of user revision) from crowdsourcing to chaining. The transferrability of these concepts can help build new chains from scratch using a broader swath of literature and past experience as guidance. Given past success in the literature survey and our case studies, we expect more approaches to tasks, objectives, strategies, and some tactics to transfer from crowdsourcing to chaining. In particular, transferring approaches to user revision and creative work from crowdsourcing could expand chaining capabilities.
5.1.1. Incorporating more user revision.
Although many chaining papers discuss involving humans into the chain hypothetically (Schick et al., 2022; Yang et al., 2022; Wu et al., 2023a; Wang et al., 2022; Wu et al., 2023b), few implement this user oversight (Park et al., 2022; Kim et al., 2023b; Mirowski et al., 2023; Zhang et al., 2023a). Our case studies showcase one example of interactions through directly manipulable outputs.
Crowdsourcing workflows integrate the user in more varied ways, including providing results without aggregation that the user can sort and filter (Lin et al., 2014a; Mohanty et al., 2019), user guidance for prototyping (Cheng and Bernstein, 2015; Kulkarni et al., 2014; Retelny et al., 2014), and worker-initiated user revision (Salehi et al., 2017). Intervention-based user revision is especially helpful for accuracy tasks. Users can break ties (Pietrowicz et al., 2013) and intervene if there are cascading errors (Kulkarni et al., 2012). Sometimes disagreements may not need to be resolved within the workflow and can instead support an explorable interface for users (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019). In creative tasks, quick iteration and choosing among uncertain options reduces the effort it takes for the user to prototype creative outputs and allows the user to guide ideation.
We recommend incorporating more approaches to user revision such as explorable interfaces, worker-initiated revision, and guided prototyping.
5.1.2. Supporting a broader range of approaches for creative work.

### v3

@@ 6.3. Case study 3: Short story generation
The final case study concerns story generation given a short text prompt. This task lies on the open-ended creativity side of the spectrum, as short stories have a wide design space of possible plots, characters, messages, imagery, and lessons. In this task, user guidance of the story involves higher level control of different plot elements, with the ability to see counterfactuals (e.g., if a story element is included or not). We implement a workflow that, compared to zero-shot approaches, enables this exploration more precisely, while also creating stories that are preferred by crowdsourced raters for imagery, originality, and style.
@@ 6.3.1. Workflow design
We adapt the Mechanical Novel (MN) (Kim et al., 2017) workflow for this task. This workflow involves three main steps (Figure 18). The first generate subtask “initializes” a story with multiple scenes based on a given story prompt (e.g., “Malcolm finds himself alone in a runaway hot air balloon and accidentally travels to a city in the sky.”). Then, the “reflect” and “revise” steps are iterated five times in a loop. In the “reflect” step, workers consider what changes might helpfully adjust the story. In the “revise” step, workers implement those changes. Each step has subtasks for parallel generation of suggestions and edits, as well as evaluate subtasks voting for the best generation.
We adapted the Crowdsourced MN to MN-LLM by incorporating fine-grained user guidance with manipulatable outputs. At the conclusion of the workflow, instead of receiving just one or more separate short stories, the user can interact with different combinations of edits to compare stories with different quality combinations and compare counterfactuals on editing decisions. To enable these outputs, we run the “revise” stage on texts of all combinations of suggestions from the “reflect” stage.
Beyond introducing control of story elements, the largest challenge was managing token context limitations. Getting response diversity with iterative improvement on texts often increases the text length (Little et al., 2010a). To account for context length, we generate stories with four scenes rather than six and iterate through scenes rather than informing workers to select scenes. In our design space this change maps to using a redundant architecture rather than a sequential architecture with a focus subtask. At the beginning of the “revise” step, we add a generation subtask that summarizes the current story to provide abbreviated context in the worker instructions. We included a hard-coded validation step with an evaluate subtask to disqualify edits that were more than 1.5 times the input text length. Unfortunately, this disqualified many highly descriptive texts and sometimes meant that no edits were made because all suggestions were too long.
@@
Our initial pilots generated editing suggestions that were very generic (e.g., “Add more detail to the story”) and were repetitive across iteration rounds. To improve the diversity of suggestion outputs, we tried changing the worker instructions to be contrastive by providing previously used “banned” suggestions. However, the outputs anecdotally seemed more likely to follow those of prior rounds after this change. We hypothesize that the length of the context and request to do two tasks in one (generate a suggestion and avoid the prior suggestions) led to a lack of diversity. Instead, we add several validation steps to promote suggestion diversity and quality. This subtask granularity worked better. In these evaluate subtasks, we prompted the LLM to ensure that 1) the suggestion did not overlap with prior suggestions or two seeded generic suggestions (“Generic recommendation to introduce conflict.” and “Generic recommendation to add detail.”), 2) the suggestion has not already been implemented in the story, and 3) that the suggestion is worded as a suggestion, not new text for the story.
6.3.2. Evaluation setup
To evaluate MN-LLM, we use the story prompts and evaluation procedure from the original paper (Kim et al., 2017) and compare MN-LLM’s results to those of two baselines.
Dataset. Our dataset consists of the five short story prompts in the Mechanical Novel paper (Kim et al., 2017).
6.3.3. Workflows enable control of plot elements
As shown in Table 3, MN-LLM generates more, longer, and higher quality suggestion combinations than the zero-shot baseline. MN-LLM does not always generate the 33 story options afforded by the initial story and combinations of 5 suggestions, as sometimes no generated suggestions passed our validation steps, or all generated fixes were too long. However, it still generates more than 20 options on average, in comparison to zero-shot’s average of 3.8.
@@ 7. Discussion
The approach of decomposing a task into a series of subtasks is conceptually the same for both crowdsourcing workflows and LLM chains. Through overlap in the literature and our case study observations, we find that the same objectives and tactics apply to both, with the exception of the choice of workers. We also find that the strategies found in crowdsourcing apply to chaining, as do their impact on objectives. Despite these similarities, LLMs and people have different abilities, strengths, and weaknesses, so we sometimes need different approaches for employing tactics to effectively implement the same strategies.
Thus, strategies and their effect on objectives transfer directly from crowdsourcing to LLMs. However, the tactic designs to support strategies can vary significantly, and could benefit from improved development and testing processes. Understanding how to effectively use tactics to implement strategies is not only a more focused goal than the connection between tactics and objectives, but it is also more generalizeable. For example, learning how to effectively get LLMs to validate their own outputs would be applicable to many different tasks and objectives.
In this section, we discuss further the similarities and differences between LLM chains and crowdsourcing workflows, lessons that exploring this intersection has taught us about using tactics to support strategies in LLM chain design, and unknowns that mark promising avenues for future discovery.
7.1. LLM stochasticity is different than human diversity
One of the key benefits of crowdsourcing workflows is that they incorporate multiple people, such that crowdworkers can fill in each other’s deficits (Huang et al., 2017; Nebeling et al., 2016; Kim and Monroy-Hernandez, 2016; Zhang et al., 2012). Therefore, requesting multiple crowdworkers for the same or sequential subtasks inherently leads to differing outputs and considerations. The diversity of perspectives happens inherently by querying different workers but is limited to available crowdworkers and thus affected by demographic and geographic disparities on crowdsourcing platforms (Difallah et al., 2018).
LLMs also have the capability to output different responses, as their training data includes information produced by many people on the Internet (Bender et al., 2021). To induce diversity, prior work has tried methods like adjusting prompts, temperature, and model types (Parameswaran et al., 2023; Park et al., 2022; Du et al., 2023; Arora et al., 2022). However, ongoing work has found that there are elements of diversity in people that are not evident in LLMs. For example, LLMs struggle to output pluralistic responses (Sorensen et al., 2024), are limited to the outputs of people with access to the Internet (Bender et al., 2021), and cannot represent identity groups (Wang et al., 2024). Our case studies suggest that LLMs need more variation in prompting and explicit intervention to get diverse responses, although we did not test this systematically. Intuitively, this makes sense because there is great diversity in the human population, but models are known to mimic the majority voice to minimize their objectives, ignoring marginalized opinions that make up a smaller percentage of the training data (Gordon et al., 2022).
Using a high temperature and prompting the LLM with a variety of different instructions can improve diversity somewhat but is unlikely to match that of crowds. Although this difference between workers may at some level seem obvious, it impacts the tactics needed to effectively use diverse responses. Diverse responses are used both to source variety and to check agreement. In crowdsourcing, the underlying tactic behind both of these goals is identical: run the same subtask with multiple workers. However, different tactics can better explicitly address these applications for LLMs.
Variety. Diverse responses are often used to support creative tasks with a variety of outputs. Consider one step in Mechanical Novel that generates suggestions for updating the story. As this is an iterative workflow, we want a variety of story update suggestions both at each iteration and across iterations. Unfortunately, even with variety in generation prompt wording, we found a low diversity of suggestion content (Table 4 “Zero-shot” column). After repeated challenges in prompt engineering to increase the variety of suggestions, we eventually intervened with an added validation step to ensure diversity of story updates. This approach was effective, but inefficient, as it filters out many generated responses. Role-based communicative architectures are also promising in giving prompts extra power to increase variety. Although giving identity based roles should treated with caution (Wang et al., 2024), giving other roles to LLMs could be a more powerful way of inducing a variety of responses (Park et al., 2022). Many roles in crowdsourcing workflows were based on organizational positions for relevant domains such as journalism (Agapie et al., 2015) or project development (Valentine et al., 2017).
To source variety, the best choice of redundant architecture can differ when using crowdworkers or LLMs. Prior crowdsourcing work has found that parallel and iterative architectures source different degrees of diversity and quality of responses in crowdsourcing depending on the task (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). For example, iterative architectures have less variation and a higher average quality of outputs, but have a lower maximum quality output (Little et al., 2010a). Some of these tradeoffs will hold with LLMs, but differences may emerge as well. For instance, some work suggests that, relative to crowdworkers who experience anchoring bias on iterative architectures, LLMs can perform divergent thinking to generate something explicitly different than provided examples (Wu et al., 2023c, a). Similar studies on LLM chains would better inform the effect of different redundant architectures on variety and quality.
Agreement. Crowdsourcing work has used agreement among crowdworkers as a proxy for confidence vs uncertainty (Liem et al., 2011; Chang et al., 2017). This agreement can be used to validate the quality of outputs, so understanding how agreement relates to quality is of critical importance. Here, we focus on agreement among outputs of identical tasks, but the difference between a different person validating the outputs of another, and a model sequentially validating its own outputs, should also be considered.
There are many potential methods for measuring confidence for LLMs, including a confidence score, checking consistency across small prompt variations, and querying multiple times for taking the majority answer (Xie et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Du et al., 2023). The effect of each of these approaches on the quality of answers is not well understood (Xie et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Du et al., 2023). It is unclear the degree to which just the consistency in one LLM’s outputs is a proxy for confidence and accuracy, versus consistency across outputs from varied prompts or models.
This more tenuous connection between agreement and correctness may have some impact on chosen tactics. For example, communicative architectures in crowdsourcing were often based on the premise that agreement among people mattered, but that may not transfer to LLMs. Consider debate-based communicative architectures. Some crowdsourcing studies only bring up debate if there’s disagreement (Chen et al., 2019; Drapeau et al., 2016; Chang et al., 2017), but LLMs may benefit more from inducing debate for all inputs, rather than if two outputs disagree.
Therefore, the differences between human diversity and LLM stochasticity impact the design of redundant and communicative architectures to support divergent thinking or agreement.
7.2. LLMs validate response quality less reliably than crowdworkers
Validation in crowdsourcing workflows often involves asking a different worker to verify if errors or desirable qualities exist in outputs. However, both the literature and our case studies show mixed results on the effectiveness of this strategy for chaining.
While some studies show that prompting a model to reflect can help achieve higher accuracy outputs (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), others caution against the same model validating its own outputs, as LLMs exhibit high confidence for incorrect prior outputs (Liang et al., 2023; Du et al., 2023). Some approaches for more robust validation include: validating with a different LLM (Liang et al., 2023), debate-based communicative structures (Liang et al., 2023; Du et al., 2023), and structural validation such as similarity across parallel redundant streams (Liem et al., 2011). One paper found that “ease of persuasion,” or how quickly an LLM relented in a debate, was more indicative of factual accuracy than other measures of confidence (Du et al., 2023).
Through our case studies, we found several helpful practices for adapting validation steps to work with LLMs.
First, validating “quality” is open-ended, so it is helpful to specify aspects of quality to check instead of using one general “quality” validation step. Crowdsourced Soylent already split the validation into separately checking for “grammar” and for “meaning” changes (Bernstein et al., 2010). However, we needed to make these checks even more targeted to adapt to LLM-specific flaws. Crowdsourced Soylent did not need to check that the rewritten text was shorter, but we added a length-checking step to Soylent-LLM after struggling to ensure a shorter length through prompting alone. We also explicitly validated that the model did not change direct quotes, a nuance that is understood in crowdworkers’ common sense. However, the subphrases from Soylent-LLM were for the most part fluent, deprioritizing the importance of the grammaticality check. In another example, crowdsourced Mechanical Novel involves one evaluate step in which crowdworkers choose the “best” suggestion. We added three validation checks to MN-LLM before this step to account for common errors in chosen suggestions. We explicitly check that LLM responses 1) do not overlap with prior suggestions, 2) are not already implemented in the story, and 3) are worded as a suggestion rather than as new text. Although we did not specify any in our case studies, LLMs may require additional validation of output structures, as text-based responses cannot be restricted by interface design the way crowdworker interfaces can.
Validation should address specific elements of quality that are aligned to worker capabilities. In our case studies, we determined what elements of quality to validate by evaluating intermediate outputs in a trial and error process. Determining these qualities can be tricky, especially as the capabilities of LLMs rapidly change. To determine what elements of quality require validation, intuition needs to be tempered by empirical testing and subtask evaluation.
Even after the set of qualities under consideration has been determined, we found LLMs could still struggle with successfully validating these qualities. To adapt to these issues, we used several approaches. First, when possible we made evaluate steps programmatic rather than LLM-based. This approach is most reliable, but there are a limited number of domains where programmatic checks are feasible. Second, we continued to break qualities down into even simpler parts when needed to align subtask difficulty with worker capabilities. Third, we simplify context to the minimum needed. Considering the story suggestion quality checks above, we found that including multiple prior suggestions, or including the entire story when not necessary, reduced the quality of validation checks.
Some types of validation remain difficult for LLMs. Anecdotally, we found that some more abstract validation tasks, such as identifying changes in meaning, were challenging to enact with LLMs. In Crowdsourced Soylent, the “Verify” step evaluates if the fix changes the meaning of the statement. This requires nuance for what elements of the meaning are necessary. Overall, supporting specificity and simplicity in validation steps were key for adapting these workflows to use LLMs.
7.3. Chains for subjective tasks can use crowdsourcing approaches to enable users to navigate the space of outputs
Some tasks have a large space of possible correct outputs, due to uncertainty or creativity, and could benefit from the user giving direct input. Under these conditions, it is helpful to draw on longstanding work on mixed-initiative user interfaces (Horvitz, 1999) to incorporate users into the workflow. Such an approach allows for parts of the workflow to be automated while others involve user input.
By enabling user initiative on workflow direction or outputs, users can navigate the space of outputs—either to explicitly take control or to learn through the process of exploration. For example, this could be helpful for determining classification label sets. Instead of predefining all possible rules in instructions, workflows can surface uncertain areas for users to make final decisions (Chang et al., 2017). In more creative ventures like prototyping, incorporating user feedback to align the output to their vision can avoid technically correct but irrelevant outputs (Kulkarni et al., 2014).
User control through prompting is brittle (Arora et al., 2022), so workflows can enable more fine grained control. LLM chains can draw from user interaction methods from crowdsourcing workflows. Several large types of user interaction emerge from the literature.
First, the user may act as a step in the workflow for validation or for clarifying context when triggered by ties (Pietrowicz et al., 2013) or crowdworkers (Kulkarni et al., 2012; Salehi et al., 2017). The main design decisions for this approach are deciding what and when the user will validate. The structure and timing of user interaction can matter, for example when supplying workers with needed context (Salehi et al., 2017).
Second, the user and crowdworkers communicate through a shared interface in a relatively unstructured cycle of interactions while creating a shared creative output. Hurdles here include making user review of outputs easier and reducing latency (Nebeling et al., 2016; Lasecki et al., 2015). Metaphorian is a chaining paper that follows a structured version of this in which users work on an interface to develop metaphors, but are limited to a few predefined actions (Kim et al., 2023b).
Third, a user may review options between iterations of the workflow (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Mirowski et al., 2023). This approach can help with exploring design spaces (Park et al., 2022; Kulkarni et al., 2014; Liu et al., 2023) and determining when the final quality threshold is met (Willett et al., 2012).
Fourth, a user may directly manipulate the output to see the breadth of user responses and to make final decisions about the output with less effort than doing the task directly (Chang et al., 2017; Chilton et al., 2013; Willett et al., 2012; Cheng and Bernstein, 2015; Lee et al., 2017; Mohanty et al., 2019; Bernstein et al., 2010). This approach enables users to surface uncertainty in the data and express control over the final output.
Many chaining papers discuss a generic possibility of a user replacing steps or intervening if needed (Schick et al., 2022; Yang et al., 2022; Wu et al., 2023a; Wang et al., 2022; Wu et al., 2023b; Zhang et al., 2024), but few actually implement this with people (Park et al., 2022; Kim et al., 2023b; Mirowski et al., 2023). We can see from crowdsourcing that there are many ways of incorporating the user that bring myriad benefits. In our case studies, we take direct manipulation from crowdsourcing and apply it to chaining, increasing the controllability of LLM outputs.
7.4. On-demand availability of LLMs reduces barriers to real-time crowdsourcing architectures
Unlike crowdworkers, LLMs are available to be queried at all times with low latency. While, in one sense, this observation is obvious, it has deep implications on tactics.
The unpredictable availability of crowdworkers presented challenges for workflows in which crowdworkers must stay available for extended periods of time or must be called on-demand (Bernstein et al., 2011). While incentive schemes could provide some real-time response, incentive tuning and implementation added significant complexity (Lasecki et al., 2013; LaToza et al., 2014; Weir et al., 2015). As a result, several crowdsourced workflows — while achieving impressive results on paper — were too complex for widespread deployment (Chen et al., 2019; Valentine et al., 2017). The availability of LLMs reduces the barriers to these workflow designs and their benefits.
Methods like retainer models that incentivize keeping crowds were important for crowdsourcing workflows that involve real-time interaction for users (Bernstein et al., 2011) and for communicative workflows like debate in which workers must remain online as other workers respond to them (Drapeau et al., 2016; Chen et al., 2019). These workflows improve latency for users and improve quality and accuracy, respectively, but the challenges of retaining workers increased barriers in crowdsourcing. Similarly, for mixed-initiative interfaces, delayed worker responses increased the latency of interactions (Lasecki et al., 2015). Reduced latency of worker responses can make user interactions more real-time.
In other workflows, such as Flash Organizations (Valentine et al., 2017), workers recruit other workers to complete subtasks that are dynamically created. Finding the right workers for the job took a median of 14 minutes with a light screening process and of 15 hours with a more intense screening process (Valentine et al., 2017). In contrast, the Orchestrator LLM of the AutoGen (Wu et al., 2023a) chain “recruits” with other calls to LLMs, which can happen flexibly and quickly (Wu et al., 2023a).
Therefore, using LLMs as workers reduces barriers to using powerful communicative workflow regimes.
7.5. LLMs need different, often simpler, subtasks
Throughout our case studies, we often needed to break crowdsourcing subtasks into multiple, more specified subtasks. Multiple factors necessitated these changes: natural language is difficult to structure (e.g., thus, we query for one item category at a time rather than eight in Cascade), LLMs have difficulty balancing multiple requirements in a subtask (Wu et al., 2023c; Madaan et al., 2023b; Hsieh et al., 2023) (e.g., in response, we split Soylent’s “Fix” step into separate edit, merge, and delete queries), and LLMs do not understand length-related constraints (e.g., to address this, we add a hard-coded validation step after a generate subtask in Mechanical Novel to check for length).
In our case studies, LLMs performed better on highly specified subtasks for chains rather than fewer, more complicated subtasks. However, this may change as newer, more powerful LLMs are introduced.
7.6. Instructions should provide the minimum necessary context
Providing global context to subtasks is widely agreed to be a critical element of subtask design and output coherence, but it can increase instruction complexity (Wu et al., 2022a; Salehi et al., 2017; Ambati et al., 2012; Luther et al., 2015). However, findings from the literature survey and case studies suggest that LLMs can respond poorly to long, complicated instructions (Du et al., 2023; Parameswaran et al., 2023; Madaan et al., 2023b; Wu et al., 2023c). This context limitation should be re-evaluated over time as LLM context windows grow.
The nuances of context inclusion appeared in our case studies. For example, we wanted to provide MN-LLM the entire short story thus far as context, as occurs in crowdsourced Mechanical Novel. To generate diverse suggestions for improvement we gave the LLM previous suggestions as context for what it should avoid. However, adding this requirement to the overall prompt anecdotally increased the output’s chance of anchoring to prior suggestions. Therefore, we made a separate evaluate subtask that determined if the new suggestion matched previous suggestions out of context from the rest of the story.
We recommend designers take into consideration that LLMs may not adequately process all the nuances of complicated instructions. Simplifying the subtask size in turn simplifies instructions and, often, the amount of necessary context.
8. Future opportunities for chaining
In this section, we highlight practices in crowdsourcing that we consider likely to be worthwhile to transfer to chaining. We encourage future work on these topics.
8.1. Creative tasks with dynamic and communicative architectures